# An Empirical Study of Training ID-Agnostic Multi-modal Sequential Recommenders

### v2

@@ 3.3. Multi-modal Fusion
As a key module that brings considerable influence in multi-modal architectures, fusion is also extensively studied in our MMSR. The fusion approaches we introduce can be divided into three classes: (1): Vanilla coarse-grained fusion methods including gate selection (Kiela et al., 2018), concatenation, summation, and FiLM (Perez et al., 2018). (2): Enhanced coarse-grained fusion methods via adding three-layer DNN to the vanilla ones, we make these modifications given that the original fusion methods may be too shallow for large text and vision encoders. (3): Fine-grained attention-based fusion approaches, including merged attention (Singh et al., 2022; Dou et al., 2022) and co-attention (Hendricks et al., 2021; Tan and Bansal, 2019). For the merged attention, linguistic and visual features are simply connected and fed to $1$ (or $n$) single-stream transformer block(s). In the co-attention module, two features are fed independently into different transformer blocks, followed by a merging block to force cross-modal interaction.
3.4. Vision-and-language Models (VLPs)
@@
The multi-modal community has developed some mature VLP models to get cross-modal semantics of text and vision information. We transfer these VLP paradigms to our MMSR to see their effect on recommendation tasks. Inspired by (Kim et al., 2021), VLPs can be split based on two points: (1) whether the two modalities have specific encoders; (2) whether the two modalities interact with each other in a deep fusion module. We first explore the impact of existing VLP architectures by directly introducing them into our framework, e.g., CLIP (Radford et al., 2021) & GroupViT (Xu et al., 2022), ViLT (Kim et al., 2021) & VisualBERT (Li et al., 2019) and FLAVA (Singh et al., 2022) (They stand for three typical paradigms of VLPs, i.e., two-stream and shallow fusion, single stream and deep fusion, two-stream and deep fusion), as is shown in Appendix E.
@@
In our experiments, we note that variations in the architecture of vision models significantly affect the overall performances. Therefore, we carry out further exploration by evolving vision representation in VLPs. We select VisualBERT (Li et al., 2019) and replace the vision input with region-based (Faster-RCNN (Faster, 2015)), grid-based (ResNet50 (He et al., 2016)), and patch-based (CLIP-ViT (Radford et al., 2021)) visual representations in Section 6.2.
3.5. Recommendation architectures
We empirically select three emblematic SR architectures to evaluate MMSR, namely Transformer-based SASRec (Kang and McAuley, 2018), GRU-based GRURec (Li et al., 2022), and CNN-based NextItNet (Yuan et al., 2019). Specifically, we juxtapose SASRec, GRURec, and NextItNet to scrutinize the interplay between the most sequential recommendation architectures and the multi-modal item encoder. In order to ensure the efficiency of the model, we default use dense next-item prediction methodologies for all SR architectures inspired by (Kang and McAuley, 2018) (see Appendix B for details).
4. Experiments
4.1. Datasets
Table 1. Dataset statistics (after preprocessing).
Dataset #users #items #cold avg.length types
@@ 4.1. Datasets
Many existing modality related SR (Hou et al., 2022b, a) primarily focus on the e-commerce domain of item images with clean and pure backgrounds (e.g., items in the Amazon dataset all belong to the ”fashion” category). However, There is insufficient evidence on whether complex multimodal information can facilitate SR with complex content (e.g. short videos with broad topics). To provide a more comprehensive evaluation, we evaluate MMSR on four real-world datasets, i.e., the HM [^1] https://www.kaggle.com/competitions/h-and-m-personalized-fashion-recommendations/overview and Amazon [^2] https://nijianmo.github.io/amazon/index.html datasets for clothing purchase, the Bili [^3] https://www.bilibili.com/ and Kwai [^4] https://www.Kwai.com/?isHome=1 dataset (Zhang et al., 2023a; Cheng et al., 2023; Li et al., 2023b) from short-video platforms. For HM, we use descriptions and covers of the products and add categorical tags to the textual features to distinguish products with the same descriptions. For Bili and Kwai datasets, the title and cover of each video are regarded as the input. We consider those with less than 10 occurrences in the training set to select cold items in the cold scenarios. More details can be seen in Table 1 and Appendix A.
4.2. Default Implementation Details
@@
4.3. Baselines
@@
Pure ID-based sequential recommenders (IDSR) assign a unique ID to each item in the dataset.
  • GRURec (Hidasi et al., 2015) employs RNN as the sequence encoder and introduces ranking loss functions for model optimization.
  • NextItNet (Yuan et al., 2019) is a CNN-based sequential recommender, which combines masked filters with dilated convolutions to increase the receptive fields.
  • SASRec (Kang and McAuley, 2018) utilizes a unidirectional Transformer as the sequence encoder, which flexibly assigns attention weights to different items in the sequence.
ID-based sequential recommenders with side features (IDSR w. side feat.) enhance pure ID-based sequential recommenders with multi-modal side features such as texts and images.
  • FDSA (Zhang et al., 2019) captures the transition patterns between items and the contextual features of items (e.g., brands, text descriptions) via feature-based self-attention blocks.
  • CARCA++ is our improved version of CARCA (Rashed et al., 2022). The original version only supports imaged-based features, and we improve it into a multi-modal version supporting both texts and images.
Modality-based sequential recommenders learn transferable item representations via item texts or images.
  • VQRec (Hou et al., 2022a) maps item texts into discrete codes and then performs the lookup operation in the code embedding table to learn transferable item representations.
  • MoRec++ is our improved version of MoRec (Yuan et al., 2023). For a fair comparison, we improve it into a multi-modal version by fusing the representations extracted from text and the vision encoders.
4.4. Loss and Negative Sampling
For consistency in our evaluation, we employ the same loss function and sampling technique for MMSR and above baseline methods. Specifically, a negative sample is randomly selected (Yuan et al., 2023) and binary cross-entropy function as the loss. For each observed interaction $<u,j>$, it can be formulated as follows:
(6) $min-\sum_{<u,i,j>\in\mathcal{S}^{sample}}\left\{\log(\sigma(\hat{y}_{ui}))+\log(1-\sigma(\hat{y}_{uj}))\right\},$
@@ 4.4. Loss and Negative Sampling
4.5. Evaluation Metrics
We adopt the standard leave-one-out strategy(Kang and McAuley, 2018) (see Appendix C for explanations) and split the three used datasets into three parts: training, validation, and testing data, and evaluate all models in terms of two Top-N ranking metrics: Hit Ratio (HR@10) and Normalized Discounted Cumulative Gain (NDCG@10, or NG@10).
5. Performance of MMSR (Q1)
To systematically explore how to build MMSR from scratch to achieve the best results without using complex tricks, we conduct extensive experiments using (1) different text and vision encoders, (2) different multimodal fusion methods, and (3) different SR architectures. The best MMSR architecture is then comprehensively compared with current baselines.
5.1. Explorations of Various Item Encoders
To explore the potential of MMSR step by step, we first try different combinations of text and vision encoders without applying existing VLP models. Concretely, we explore the potential of our MMSR in unimodal mode [^6] Note that MMSR-T and MMSR-V in the tables are the ablation versions of our MMSR with only the text and vision information., then we construct the complete multi-modal MMSR by merging both TE and VE, i.e., we set RoBERTa (Liu et al., 2019) as the default TE and combine it with various VE, and choose MAE (He et al., 2022) as the default VE to integrate with other TE in reverse. For all of them, we select a one-block merge attention module as the fusion approach. All explorations are evaluated over HM and Bili datasets with SASRec, and the results are reported in Table 2.
Table 2. Comparisons of MMSR with different TE(i.e., BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019) and ELECTRA (Clark et al., 2020)) and VE (i.e., ResNet50 (He et al., 2016), MAE B-224/16 (He et al., 2022), and Swin B-224/16 (Liu et al., 2021)). All TE are in the base size. RoBERTa and MAE are default TE and VE, respectively.
@@ 5.1. Explorations of Various Item Encoders
Various Text Encoders. There is no significant difference in performance for different TE in both MMSR-T and MMSR, while RoBERTa(Liu et al., 2019) always seems to achieve the best performance. This suggests that the capabilities of various TEs in feature representation have become very close, indicating that text representation techniques have matured considerably.
Various Vision Encoders. MAE (He et al., 2022) or Swin (Liu et al., 2021) perform the best in MMSR. However, the results of MAE (He et al., 2022) and ResNet (He et al., 2016)  [^7] Like PixelBERT (Huang et al., 2020), ResNet is used as the backbone to reduce the dimension of pixel features and we selected 49 vectors from each image. are exactly opposite. For the poor performance of ResNet (He et al., 2016) in MMSR, we guess it is because the current fusion approach cannot digest the outputs of ResNet (He et al., 2016) (CNN-based) and RoBERTa (Liu et al., 2019) (transformer-based) well. For MAE (He et al., 2022), the default high mask ratio (75%) may be the reason for its poor performance in MMSR-V.
Results. The performance of MMSR-T is generally better than that of MMSR-V, which may be that the dense semantics of textual information is easier to learn in SR (Hou et al., 2022b). The fully transformer-based (e.g., RoBERTa and MAE or Swin) perform the best in MMSR, which suggests the superiority of unified architectures.
Table 3. Comparison MMSR with various fusion methods. Coarse: MMSR with coarse fusion strategies. Coarse+: add a three-layer DNN based on the coarse fusion. CoAtten-1: co-attention fusion of 1 layer. MergeAttn-$n$: merge-attention fusion of $n$ layer.
@@
Table 4. Comparison of IDSR and MMSR with various SR architectures. HR@10 is as the default metric here.
@@
Table 5. Performance comparison (%) on four datasets. The best and the second-best performance in each row are bolded and underlined, respectively. Improvements compared with the best baseline method are indicated in the last column.
@@
5.2. Impact of Various Fusion Approaches
In this section, we show investigations on various multi-modal fusion strategies for MMSR with a default of TE, VE and RS architecture (i.e., RoBERTa (Liu et al., 2019), MAE (He et al., 2022), and SASRec (Kang and McAuley, 2018)). Specifically, we explore four vanilla coarse-grained fusion methods (i.e., gate selection (Kiela et al., 2018), concatenation, summation, and FiLM (Perez et al., 2018)), four enhanced coarse-grained fusion methods (with a three-layer DNN applied to the previous ones), and two fine-grained fusion methods (i.e., co-attention (Hendricks et al., 2021) and merge-attention (Dou et al., 2022)).
Coarse-grained fusion. As shown in Table 3, MMSR with coarse-grained fusion lacks competitiveness, especially in the HM dataset. Among them, gate (Kiela et al., 2018) and FiLM (Perez et al., 2018) do slight favor to MMSR while summation and concatenation seriously deteriorate the performance. Comparing Coarse with Coarse++, we notice that added DNN drives all the shallow fusions to similar performance. It seems the three-layer DNN overwhelms the multi-modal fusion.
Fine-grained fusion. The results in Table 3 show that merge-attention outperforms co-attention. We also find that there is no improvement even though we deepen the merge-attention from one to four transformer blocks, and the results on the Bili dataset are even inversely proportional to the depth. It may be that the amount of data is not enough to support the learning of deep fusion. It suggests that VLP models should be considered if we want to use deeper fusion modules (see Section 6.1).
@@ 5.2. Impact of Various Fusion Approaches
5.3. MMSR with Various SR Architectures
Based on the above explorations, we choose powerful RoBERTa (Liu et al., 2019) and Swin (Liu et al., 2021) as text and vision encoders, respectively. To validate the robustness of our MMSR, we consider the effect of various architectures, including SASRec (Kang and McAuley, 2018), GRURec (Hidasi et al., 2015), and NextItNet (Wang et al., 2022), as shown in Figure 2 (right). We evaluate the performance over four recommendation datasets (i.e., HM, Bili, Kwai, and Amazon).
Results. As shown in Table 4, MMSR configured with SASRec shows strong competitiveness. Furthermore, no matter which SR architecture is given, the performance of traditional ID-based SR (IDSR) are much lower than that of MMSR. The ablation study shows that single-modal information, is still very competitive, even surpassing IDSR in some cases. This suggests that current CV and NLP techniques have enabled id-agnostic recommendation models to excel, highlighting the potential of the MMSR paradigm.
5.4. Overall Performance Comparisons
We evaluate MMSR and other baseline methods on 4 source datasets. From the results of Table 5, we can reach the following conclusions:

### v5

@@ 3.3 Multi-modal Fusion
As a key module that brings considerable influence in multi-modal architectures, fusion is also extensively studied in our MMSR. The fusion approaches we introduce can be divided into three classes: (1): Vanilla coarse-grained fusion methods including gate selection [46], concatenation, summation, and FiLM [47]. (2): Enhanced coarse-grained fusion methods via adding three-layer DNN to the vanilla ones, we make these modifications given that the original fusion methods may be too shallow for large text and vision encoders. (3): Fine-grained attention-based fusion approaches, including merged attention [30, 48] and co-attention [49, 50]. For the merged attention, linguistic and visual features are simply connected and fed to $1$ (or $n$) single-stream transformer block(s). In the co-attention module, two features are fed independently into different transformer blocks, followed by a merging block to force cross-modal interaction.
3.4 Vision-and-language Models (VLPs)
@@
The multi-modal community has developed some mature VLP models to get cross-modal semantics of text and vision information. We transfer these VLP paradigms to our MMSR to see their effect on recommendation tasks. Inspired by [28], VLPs can be split based on two points: (1) whether the two modalities have specific encoders; (2) whether the two modalities interact with each other in a deep fusion module. We first explore the impact of existing VLP architectures by directly introducing them into our framework, e.g., CLIP [26] & GroupViT [27], ViLT [28] & VisualBERT [29] and FLAVA [30] (They stand for three typical paradigms of VLPs, i.e., two-stream and shallow fusion, single stream and deep fusion, two-stream and deep fusion).
@@
In our experiments, we note that variations in the architecture of vision models significantly affect the overall performances. Therefore, we carry out further exploration by evolving vision representation in VLPs. We select VisualBERT [29] and replace the vision input with region-based (Faster-RCNN [31]), grid-based (ResNet50 [23]), and patch-based (CLIP-ViT [26]) visual representations in Section 6.2.
3.5 Recommendation Architectures
To evaluate the MMSR framework, we systematically selected three representative SR architectures: Transformer-based SASRec [3], GRU-based GRURec [51], and CNN-based NextItNet [2]. We analyze the synergy between these popular sequential recommendation architectures and multi-modal item encoders, focusing on their ability to integrate and leverage multi-modal data effectively. We use joint distributions to capture comprehensive sequence relationships, as shown in the equation:
$R(x)=\prod\limits_{i=0}^{l}r(x_{i}|x_{0:l-1},\theta)r(x_{0})$ (3)
where, $r(x_{i}|x_{0:l-1},\theta)$ represents the conditional probability of the $i$-th item, given all preceding items $x_{0:i-1}$.
We experiment with various hyperparameters: the number of layers is set to 2, 4, and 2 for SASRec, GRURec, and NextItNet, respectively, after experimentation with options in $\{1,2,4,8\}$. For SASRec, we configure the attention heads to 2, aiming to optimize the interaction between sequential processing and multi-modal integration for enhanced recommendation performance.
4 Experiments
4.1 Datasets
Many existing modalitiey-related SR [17, 18] primarily focus on the e-commerce domain of item images with clean and pure backgrounds (e.g., items in the Amazon dataset all belong to the ”fashion” category). However, There is insufficient evidence on whether complex multimodal information can facilitate SR with complex (e.g. short videos with broad topics).content To provide a more comprehensive evaluation, we evaluate MMSR on four real-world datasets, i.e., the HM [^1] https://www.kaggle.com/competitions/h-and-m-personalized-fashion-recommendations/overview and Amazon [^2] https://nijianmo.github.io/amazon/index.html datasets for clothing purchase, the Bili [^3] https://www.bilibili.com/ and Kwai [^4] https://www.Kwai.com/?isHome=1 dataset from short-video platforms. For HM, we use descriptions and covers of the products and add categorical tags to the textual features to distinguish products with the same descriptions. For Bili and Kwai datasets, the title and cover of each video are regarded as the modality input. We consider those with less than 10 occurrences in the training set to select cold items in the cold scenarios.
The HM and Amazon datasets include customer-product (user-item) interactions and customer and product metadata, e.g., category and color, descriptions, and covers. In this paper, we only use texts and images, adding some category information as part of the description to further distinguish different items. Following [2], we removed users with less than 5 interactions simply, and randomly selected 200K and 100K users.
To build the Bili dataset, we randomly crawl videos (less than 10 minutes in duration) of different categories from the online platform. We record the public comment information of these videos as interactions (i.e., we think the user would be interested in this video if a user comments on a video). We do not capture any user’s personal information about privacy, whose IDs are already anonymous after data masking. The video metadata such as cover and title are recorded. The Kwai dataset is collected similarly to the Bili dataset. Then we also removed users with less than 5 interactions and randomly selected about 100K and 200K users from the Bili and Kwai datasets, respectively.
Besides, for all datasets, we consider cold scenarios (different from regular scenarios). That is, we count the interactions of all items in the training set and consider those items with less than 10 occurrences as cold items [8]. Then, we truncate the complete user sequence to get some sub-sequences with a cold item at the end, all of which are used to evaluate the capability of MMSR in the cold scenario. The preprocessed dataset statistics are presented in Table  I.
TABLE I: Dataset statistics (after preprocessing). Specifically, users, items, and cold denote the numbers of users, items, and truncated users sub-sequences with a cold item at the end, respectively. The avg. denotes the average sequence length of users’ historical interactions.
Dataset #users #items #cold avg. types
@@ 4.1 Datasets
4.2 Default Implementation Details
@@
[…]
Following [52], we rank the ground-truth target item by comparing it with all the other items in the item pool. Finally, we report the results of the testing set by choosing checkpoints with the best validation performance.
4.3 Baselines
@@
Pure ID-based sequential recommenders (IDSR) assign a unique ID to each item in the dataset. Item IDs are subsequently converted into embedding vectors and then fed into a sequence encoder to extract user preferences. However, as item IDs are not shareable across domains and platforms, these methods are difficult to transfer to new recommendation scenarios. Typical examples are listed as follows:
  • GRURec [1] employs RNN as the sequence encoder and introduces ranking loss functions for model optimization.
  • NextItNet [2] is a CNN-based sequential recommender, which combines masked filters with dilated convolutions to increase the receptive fields.
  • SASRec [3] utilizes a unidirectional Transformer as the sequence encoder, which flexibly assigns attention weights to different items in the sequence.
ID-based sequential recommenders with side features (IDSR w. side feat.) enhance pure ID-based sequential recommenders with multi-modal side features such as texts and images. The core idea is to integrate multi-modality contents into the backbone of ID-based sequence encoders in order to improve recommendation performance. Such designing is still constrained under the ID-based paradigm, thus still falls short of transferability. Typical examples are listed as follows:
  • FDSA [11] captures the transition patterns between items and the contextual features of items (e.g., brands, text descriptions) via feature-based self-attention blocks.
  • CARCA++ is our improved version of CARCA [15]. It utilizes cross-attention to capture the correlation between item IDs and contextual multi-modal features. The original version only supports imaged-based features, and we improve it into a multi-modal version.
Modality-based sequential recommenders learn transferable item representations via item texts or images. Typical examples are listed as follows:
  • VQRec [18] maps item texts into discrete codes and then performs the lookup operation in the code embedding table to learn transferable item representations.
  • MoRec++ is our improved version of MoRec [8]. For a fair comparison, we improve it into a multi-modal version by fusing the representations extracted from text and the vision encoders.
4.4 Loss and Negative Sampling
We adopt the same negative sampling strategy  [^6] Some literature [53, 54] shows other advanced negative sampling strategies, which does not affect the fairness of our work. in [8] which is absolutely fair to MMSR and other SR methods. For consistency in our evaluation, we employ the same loss function and sampling technique for MMSR and above baseline methods. Specifically, a negative sample is randomly selected [8] and binary cross-entropy function as the loss. For each observed interaction $<u,j>$, it can be formulated as follows:
$min-\sum_{<u,i,j>\in\mathcal{S}^{sample}}\left\{\log(\sigma(\hat{y}_{ui}))+\log(1-\sigma(\hat{y}_{uj}))\right\},$ (6)
@@ 4.4 Loss and Negative Sampling
4.5 Evaluation Metrics
We adopt the standard leave-one-out strategy [3] and split the three used datasets into three parts: training, validation, and testing data, and evaluate all models in terms of two Top-N ranking metrics: Hit Ratio (HR@10) and Normalized Discounted Cumulative Gain (NDCG@10, or NG@10 for short).
5 Performance of MMSR (Q1)
To systematically explore how to build a general MMSR framework to achieve the best results, we conduct extensive experiments using (1) different text and vision encoders, (2) different multimodal fusion methods, and (3) different SR architectures. The best MMSR architecture is then comprehensively compared with current baselines.
5.1 Explorations of Various Item Encoders
To explore the potential of MMSR step by step, we first try different combinations of text and vision encoders without applying existing VLP models. Concretely, we explore the potential of our MMSR in unimodal mode [^7] Note that MMSR-T and MMSR-V in the tables are the ablation versions of our MMSR with only the text and vision information., then we construct the complete multi-modal MMSR by merging both TE and VE, i.e., we set RoBERTa [21] as the default TE and combine it with various VE, and choose MAE [25] as the default VE to integrate with other TE in reverse. For all of them, we select a one-block merge attention module as the fusion approach. All explorations are evaluated over HM and Bili datasets with SASRec, and the results are reported in Table II.
TABLE II: Comparisons of MMSR with different TE(i.e., BERT [6], RoBERTa [21] and ELECTRA [22]) and VE (i.e., ResNet50 [23], MAE B-224/16 [25], and Swin B-224/16 [45]). All TE are in the base size. RoBERTa and MAE are default TE and VE, respectively.
@@ 5.1 Explorations of Various Item Encoders
Various Text Encoders. We do not observe significant differences in performance for different TE in both MMSR-T and MMSR. This suggests that the capabilities of various TEs in feature representation have become very close, indicating that text representation techniques have matured considerably.
Various Vision Encoders. MAE [25] or Swin [45] perform the best in MMSR. However, the results of MAE [25] and ResNet [23]  [^8] Like PixelBERT [24], ResNet is used as the backbone to reduce the dimension of pixel features and we selected 49 vectors from each image. are exactly opposite. For the poor performance of ResNet [23] in MMSR, we guess it is because the current fusion approach cannot digest the outputs of ResNet [23] (CNN-based) and RoBERTa [21] (transformer-based) well. For MAE [25], the default high mask ratio (75%) may be the reason for its poor performance in MMSR-V.
Results. The performance of MMSR-T is generally better than that of MMSR-V, which may be that the dense semantics of textual information is easier to learn in SR [17]. The fully transformer-based (e.g., RoBERTa and MAE or Swin) perform the best in MMSR, which suggests the superiority of unified architectures. The MMSR with a comprehensive combination of text and vision encoders demonstrates its superiority.
TABLE III: Comparison of MMSR with various fusion methods. RoBERTa (base) [21] and MAE B-224/14 [25]) are used as text and vision encoders, respectively. Coarse: MMSR with coarse fusion strategies. Coarse+: add a three-layer DNN based on the coarse fusion. CoAtten-1: co-attention fusion of 1 layer. MergeAttn-$n$: merge-attention fusion of $n$ layer.
@@
TABLE IV: Comparison of IDSR and MMSR with various SR architectures. HR@10 is as the default metric here. RoBERTa (base)[21] and Swin B-224/14 [45]) are selected as the text and vision encoders. The best results are bolded. ‘Improv.’ is the relative improvement of the MMSR compared to the best IDSR.
@@
TABLE V: Performance comparison (%) on four datasets. The best and the second-best performance in each row are bolded and underlined, respectively. Improvements compared with the best baseline method are indicated in the last column.
@@
5.2 Impact of Various Fusion Approaches
In this section, we show investigations on various multi-modal fusion strategies for MMSR with a default of TE, VE and RS architecture (i.e., RoBERTa [21], MAE [25], and SASRec [3]). Specifically, we explore four vanilla coarse-grained fusion methods (i.e., gate selection [46], concatenation, summation, and FiLM [47]), four enhanced coarse-grained fusion methods (with a three-layer DNN applied to the previous ones), and two fine-grained fusion methods (i.e., co-attention [49] and merge-attention [48]).
Coarse-grained fusion. As shown in Table III, MMSR with coarse-grained fusion lacks competitiveness, especially in the HM dataset. Among them, gate [46] and FiLM [47] do slight favor to MMSR while summation and concatenation seriously deteriorate the performance. Comparing Coarse with Coarse++, we notice that added DNN drives all the shallow fusions to similar performance. It seems the three-layer DNN overwhelms the multi-modal fusion.
Fine-grained fusion. The results in Table III show that merge-attention outperforms co-attention. We also find that there is no improvement even though we deepen the merge-attention from one to four transformer blocks, and the results on the Bili dataset are even inversely proportional to the depth. It may be that the amount of data is not enough to support the learning of deep fusion. It suggests that VLP models should be considered if we want to use deeper fusion modules (see Section 6.1).
@@ 5.2 Impact of Various Fusion Approaches
5.3 MMSR with Various SR Architectures
Based on the above explorations, we choose powerful RoBERTa [21] and Swin [45] as text and vision encoders, respectively. To validate the robustness of our MMSR, we consider the effect of various architectures, including SASRec [3], GRURec [1], and NextItNet [19], as shown in Figure 2 (right). We evaluate the performance over four recommendation datasets (i.e., HM, Bili, Kwai, and Amazon).
Results. As shown in Table IV, MMSR configured with SASRec shows strong competitiveness. Furthermore, no matter which SR architecture is given, the performance of traditional ID-based SR (IDSR) are much lower than that of MMSR. The ablation study shows that single-modal information, is still very competitive, even surpassing IDSR in some cases. This suggests that current CV and NLP techniques have enabled id-agnostic recommendation models to excel, highlighting the potential of the MMSR paradigm.
5.4 Overall Performance Comparisons
We evaluate MMSR and other baseline methods on 4 source datasets. From the results of Table V, we can reach the following conclusions: